OpenStudy (susanna):

Find the values of x for which the equation cos x=-1 is true. let k represent an integer?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \cos(x)=-1 }\) If you look at the unit circle, you will likely know that \(\color{#000000 }{ \displaystyle \cos(\pi)=-1 }\) Then, other coterminal foll from this, \(\color{#000000 }{ \displaystyle \cos(\pi)= \cos(3\pi)= \cos(5\pi) ...=-1 }\)

OpenStudy (mathstudent55):